OpenStudy (anonymous):

Find the monic polynomial, p(x) of degree 3 which has a remainder of -1 when divided by x-1, a remainder of 1 when divided by x+1 and p(0) = 1

OpenStudy (anonymous):

\[p(x)=ax^3+bx^2+cx+1\]

OpenStudy (anonymous):

yeah... i know that but i cant do the rest

OpenStudy (anonymous):

long division?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

since If a polynomial p(x) is divided by (x-r), then the remainder is a constant given by p(r) . \[p(1)=-1,p(-1)=1\]

OpenStudy (anonymous):

\[a+b+c+1=-1\] \[-a+b-c+1=1\]